OpenStudy (anonymous):

The slope of y = x 2 + kx - 1 at x = 1 is 5. Find k. 0 2 3 5 @Data_LG2 @FibonacciChick666 is it 3?

OpenStudy (fibonaccichick666):

work?

OpenStudy (anonymous):

i used the nth rule to find derivative: y'=2x+k-0 or y'=2x+k. Since y'=5 and x=1, i solved the following 5=2*1+k-------> k=3

OpenStudy (fibonaccichick666):

yep yep I agreee

OpenStudy (anonymous):

thank you very much. now i at least can be sure i know something about derivatives!!

OpenStudy (fibonaccichick666):

:) no probs

OpenStudy (anonymous):

keep up the good work @ksanka ! :)